You are trying to estimate the cost of capital to use in assessing a new investment...

  1. You are trying to estimate the cost of capital to use in assessing a new investment venture in the entertainment business, for TechSmith Inc, a publicly traded electronics company. You have been provided the following information:
  • The beta for TechSmith, based upon stock prices for the last 5 years, is 1.80 but the unlevered beta for entertainment companies is 1.20.
  • At the moment, TechSmith has one bank loan outstanding, with a principal payment due of $250 million at the end of 10 years and interest payments of $ 10 million every year for the next 10 years. TechSmith also has lease commitments of $20 million a year, due at the end of each year for the next 8 years.
  • TechSmith has 60 million shares outstanding, with a stock price of $20/share.
  • TechSmith has a bond rating of BB, with a default spread of 4.5% over the risk-free rate. The marginal tax rate if 40% but TechSmith’s effective tax rate is25%.
  • The risk-free rate is 3% and the equity risk premium is 6%.

  1. Estimate the market value of the interest-bearing debt.
  2. Estimate the debt value of lease commitments.
  3. Estimate the cost of capital for this project, assuming that it will be funded using the same debt ratio that TechSmith uses to fund itself today.
Answer #1

1. Market Value of interest-bearing debt is calculated as below

C[(1 – (1/((1 + Kd)^t)))/Kd] + [FV/((1 + Kd)^t)]

Where C is the interest expense (in dollars)

Kd is the current cost of Debt (in percentages)

T is the weighted average maturity (in years)

FV represents the total debt

Principal payment 250$ and interest 10$ every year

cost of capital using camp formulae rf+beta(rm-rf) = 3%+6% = 9%

= (10,000,000*(1-1/1.09^10)/0.9) = 6,42,222

add the above amount to

principal calculation i.e 250,000,000/1.09^10) = 105,602,701

and the total market value is = 106244923 is the estimated hypothetical value of the bond to arrive at market debt value of company is 356,244,923

Part 2 :

Debt value of lease commitments i.e 20 milllion$ each ear for next 8 years